OpenStudy (anonymous):

rewrite the expression sin^2(2x) in terms of the first power of the cosine

OpenStudy (anonymous):

\[1- \cos4x \div 2 \]

OpenStudy (anonymous):

Thanks! Can you provide an explanation to the solution?

OpenStudy (raden):

use the identity : sin^2 (p) = (1 - cos(2p))/2 now, change p = 2x, thus, sin^2 (2x) = (1 - cos(2 * 2x))/2 = (1 - cos(4x))/2

OpenStudy (anonymous):

Nice! Great explanation! Thanks so much! ;)

OpenStudy (raden):

you;re welc0me